A Roman statue is 3 times as old as a Florentine statue. One hundred years from now the roman statue will be twice as old. How old is the roman statue?
Answer by jorel1380(3719)

R=3F
R+100=2(F+100)
3F+100=2F+200
F=100
R=300
The Roman statue is 300 years old. The Florentine statue is 100 years old.